# Reinforcement Learning based Autonomous Multi-Rotor Landing on Moving Platforms

### v2

@@ 1 Introduction
Section 2 presents related work, followed by Section 3 describing the proposed approach in detail. Section 4 presents the implementation and experimental setup. The results are discussed in Section 5, with comments on future work given in Section  6.
@@ 3.3.1 Dynamic Decoupling
The nonlinear dynamics of a multi-rotor vehicle such as a quadcopter can be greatly simplified when the following assumptions are applied. i) Linearization around hover flight, ii) small angles iii) a rigid, symmetric vehicle body, iv) no aerodynamic effects [26]. Under these conditions, the axes of motion are decoupled. Thus, four individual controllers can be used to independently control the movement in longitudinal, lateral, vertical direction and around the vertical axis. For this purpose, low-level controllers track a setpoint $\theta_{ref}$ for the pitch angle (longitudinal motion), $\phi_{ref}$ for the roll angle (lateral motion), $T_{ref}$ for the total thrust (vertical motion) and $\psi_{rel}$ for the yaw angle (rotation around vertical axis). We leverage this fact in our controller structure as is further illustrated in Section 4. Furthermore, this enables us to introduce a learning task for longitudinal 1D motion only. Its purpose is to learn a policy for the pitch angle setpoint $\theta_{ref}$ that induces a 1D movement in longitudinal direction, allowing the vehicle to center over the platform moving in longitudinal direction as well. After the training, we then apply a second instance of the trained agent for controlling the lateral motion, where it produces set point values for the roll angle $\phi_{ref}$. This way, full 2D motion capability is achieved. We use the basis learning task to compose a curriculum later in Section 3.4. PID controllers ensure that $\phi_{rel}=0$\mathrm{r}\mathrm{a}\mathrm{d}$,\psi_{rel}=0$\mathrm{r}\mathrm{a}\mathrm{d}$$ and $v_{z_{rel}}=\text{const}<0$\mathrm{m}\mathrm{/}\mathrm{s}$$ during training.
@@ 3.3.4 Discrete State Space
where $p_{max},v_{max},a_{max}$ are the values used for the normalization of the observations. The reason for the clipping is that our state space discretization technique, a crucial part of the sequential curriculum, is derived assuming a worst case scenario for the platform movement in which the multi-rotor vehicle is hovering (see Section 3.4) and where scaling an observation with its maximum values, $p_{max},v_{max}$ and $a_{max}$, would constitute a normalization. However, once the multi-rotor vehicle starts moving too, the scaled observation values could exceed the value range of $[-1,1]$. Clipping allows the application of the discretization technique also for a moving multi-rotor vehicle. Next, we define a general discretization function $d(x,x_{1},x_{2})$ that can be used to map a continuous observation of the environment to a discrete state value.
@@ 3.3.5 Goal State
$s^{*}=\left\{1,1,1,*\right\}.$ (18)
This means the goal state is reached if $-p_{goal}\leq p_{x}\leq p_{goal}$, $-v_{goal}\leq v_{x}\leq v_{goal}$ and $-a_{goal}\leq a_{x}\leq a_{goal}$ regardless of the value of $i_{\theta}$.
@@ 3.3.6 Reward Function
The weights are negative so that decreasing the relative position, relative velocity or the pitch angle yields positive reward values. Clipping $r_{p}$ and $r_{v}$ to ±$\pm r_{p,max}$ and ±$\pm r_{v,max}$, as well as scaling $r_{\theta}$ and $r_{dur}$ with $v_{lim}$ is necessary due to their influence on the value of $r_{max}$. $r_{max}$ denotes the maximum achievable reward in a non-terminal timestep if $v_{x}\leq v_{lim}$ and $a_{x}\leq a_{lim}$ and thus complies with the limits set by the motion scenario described in Section 3.4 during the derivation of the sequential curriculum. To ensure the applicability of the curriculum also in situations in which that compliance is not given is why the clipping is applied in (20) and (21). $r_{max}$ plays a role in the scaling of Q-values that is performed as part of the knowledge transfer during different curriculum steps. It is defined as follows
@@ 3.4 Curriculum and Discretization
In our novel approach, we introduce a similar multiresolution technique combined with transfer learning as part of a sequential curriculum [17] to accelerate and stabilize training even further. Each new round of training in the multiresolution setting constitutes a different curriculum step. A curriculum step is an instance of the basis learning task introduced in Section 3.3. This means in particular that each curriculum step has its own Q-table. Since the size of the state and action space is the same throughout all steps, knowledge can be easily transferred to the subsequent curriculum step by copying the Q-table that serves then as a starting point for its training. Fig. 3 illustrates the procedure of the sequential curriculum.
@@
Furthermore, the discretization introduced in Section 3.3.4, which is now associated with a curriculum step, covers only a small part of the continuous observation space with few discrete states. In order to completely capture the predominant environment characteristics of the subspace in which the agent is expected to learn the task, it is necessary for it to have the knowledge of the state space topology [1]. Against the background of the multiresolution approach, this can be rephrased as the question of how to define suitable goal regions, i.e., the size of the goal states. To this end, we again leverage the insight that accurate knowledge about the relative position and velocity is only required when landing is imminent. We begin by choosing an exponential contraction of the size of the unnormalized goal state $p_{c,x,goal,i}$ associated with the $i^{th}$ curriculum step such that
@@
where $10<\sigma<$ 1 is a contraction factor, $l_{mp}\in\mathbb{R}[$\mathrm{m}$]$ is the edge length of the squared moving platform, $x_{max}$ is the boundary of the fly zone $\mathcal{F}$ (see Section 3.1.2) and $n_{cs}$ is the number of curriculum steps following the initial step. The contraction factor $\sigma$ plays a role in how easy it is for the agent to reach the goal state. If it is set too low, the agent will receive success rewards only rarely, thus leading to an increased training time. (34) can be solved for $n_{cs}$ because $\sigma$, $l_{mp}$ and $x_{max}$ are known parameters. To determine the relationship between the contraction of the positional goal state and the contraction of the velocity goal state, we need a kinematics model that relates these variables. For this purpose, we envision the following worst case scenario for the relative motion (1). Assume that the vehicle hovers right above the platform which is located on the ground in the center of the fly zone $\mathcal{F}$, with $v_{x}=0$\mathrm{m}\mathrm{/}\mathrm{s}$$ and $\psi_{rel}=0$. Now, the platform starts to constantly accelerate with the maximum possible acceleration $a_{mp,max}$ defined by (7) until it has reached $x_{max}$ after a time $t_{0}=\sqrt{2x_{max}/a_{mp,max}}$. This is considered to be the worst case since the platform never slows down, unlike the rectilinear periodic movement (6) used for training. The evolution of the continuous observations over time is then expressed by
@@
which constitute a simple kinematics model that relates the relative position and relative velocity via the time $t$. Next, we consider $p_{max}=x_{max}$, $v_{max}=a_{mp,max}t_{0}$ and $a_{max}=a_{mp,max}$ for the normalization (11) - (13) and discretize (35) - (37) via the time $t_{i}=\sigma^{i}t_{0}$. This allows us to determine the values of $p_{lim,i}$, $v_{lim,i}$, $a_{lim,i}$ and $p_{goal,i}$, $v_{goal,i}$, $a_{goal,i}$. They are required for the basis learning task presented in Section 3.3 that now constitutes the $i^{th}$ curriculum step. For this purpose, we define with $i\in\left\{0,1,\ldots,n_{cs}\right\}$
@@
During the training of the different curriculum steps the following episodic terminal criteria have been applied. On the one hand, the episode terminates with success and the success reward $r_{suc}$ is received if the goal state $s^{*}$ of the latest curriculum step is reached if the agent has been in that curriculum step’s discrete states for at least one second without interruption. This is different to all previous curriculum steps where the success reward $r_{suc}$ is received immediately after reaching the goal state of the respective curriculum step as is explained in Section 3.3.6. On the other hand, the episode is terminated with a failure if the multi-rotor vehicle leaves the fly-zone $\mathcal{F}$ or if the success terminal criterion has not been met after the maximum episode duration $t_{max}$. The reward received by the agent in the terminal time step for successful or failing termination of the episode is defined by (29). Once the trained agent is deployed, the agent selects the actions based on the Q-table of the latest curriculum step to which the continuous observations can be mapped (see Fig. 3).
@@ 3.5 Hyperparameter Determination
We leverage the discrete action space (8) to determine the hyperparameters agent frequency $f_{ag}$ and maximum pitch angle $\theta_{max}$ in an interpretable way. The purpose is to ensure sufficient maneuverability of the UAV to enable it to follow the platform. For sufficient maneuverability, the UAV needs to possess two core abilities. i) Produce a maximum acceleration bigger than the one the platform is capable of and ii) change direction of acceleration quicker than the moving platform. Against the background of the assumptions explained in Section 3.3.1 complemented with thrust compensation, we consider the first aspect by the maximum pitch angle $\theta_{max}$
@@ 4.7 Cascaded PI Controller Baseline
We apply separate instances for controlling the longitudinal, lateral and vertical movement of the UAV. The properties of the different instances are summarized in Table 4. The outer cascade tracks the setpoint $0$\mathrm{m}$$ for the relative position ${}_{s}r_{rel}$ between the multi-rotor vehicle and the moving platform. For this purpose, it generates a setpoint for the relative velocity ${}_{s}\dot{r}_{rel}$ that is tracked by the inner cascade. The inner cascade aims to reach the velocity setpoint by computing setpoints for the attitude angle $\phi_{ref},\theta_{ref}$ or the thrust $T_{ref}$, respectively. The attitude angle or thrust is then tracked by the multi-rotor vehicle’s low level controllers. In order to achieve comparability between our RL approach and the cascaded PI controller, we limit the velocity setpoints generated by the outer cascade to the value of $v_{max}=a_{mp,max}t_{0}$ (see Section 3.4) and the attitude angles generated by the inner cascade to the same angles determined using (44) for the training scenario associated with the agent chosen for comparison.
@@ 3 $9,7,8,8$ $8,6,7,7$ $7,6,8,8$ $6,7,6,8$
They were determined over a larger set of RPMs than in the baseline method and indicate a good performance of the approach. The success rates become higher when the platform velocity is lower during evaluation compared to the one applied during training. This is to be expected since the equations presented in Sec. 3.5 for hyperparameter determination ensure a sufficient maneuverability up to the velocity of the rectilinear periodic movement used in training. For the training case RPM 0.8, the fourth agent has a comparably poor performance. For the static platform the agent occasionally suffers from the problem of motion initiation as described in Section 4.6, despite setting $\psi_{rel}=\pi/4$\mathrm{r}\mathrm{a}\mathrm{d}$$. For the evaluation case where the platform moves with RPM 0.4, the reason is an oscillating movement, which causes the agent to occasionally overshoot the platform. A similar problem arises for agent four of training case hardware where the agent occasionally expects a maneuver of the platform during the landing procedure. As a consequence, this agent achieves a success rate of only $82\%$ for a static platform. However, for higher platform velocities, the success rate improves significantly.
@@ 5.2 Selection of Agents for further Evaluation
We compare its results with the RL baseline method and the cascaded PI controller in Table 8. The comparison shows that our approach is able to outperform the RL baseline method. For the RPM $0.4$ evaluation scenario with noiseless observations we achieve a success rate of $99\%$, which is $+8\%$ better than the baseline. For the RPM $1.2$ evaluation scenario, our method is successful in $99\%$ of the landing trials, increasing the RL baseline’s success rate by $+26\%$. Note that the maximum radius of the rectilinear periodic movement is $r_{mp}=\sim 2.5$\mathrm{m}$$ in the RL baseline and $r_{mp}=2$\mathrm{m}$$ in our approach. However, the value used for our approach poses a more difficult challenge since the acceleration acting on the platform is higher, due to the same maximum platform velocity, see Table 2. Furthermore, our method requires $\sim 80\%$ less time to train and $53\%$ less episodes. Comparison with the the cascaded PI controller shows that our method is capable of similar performance than the cascaded PI controller in terms of success rates achieved in the two comparison scenarios. In order to achieve comparability in the first place, we limited the control efforts of the longitudinal and lateral controller’s outer and inner cascade (see Table 4) to values also derived for the training case RPM 1.6. We obtain $u_{o}=v_{max}=a_{mp,max}t_{0}=3.39$\mathrm{m}\mathrm{/}\mathrm{s}$$ as described in Section 3.4 and $u_{i}=\theta_{ref}=\phi_{ref}=21.38^{\circ}$ by means of (44), respectively. As is further shown by Table 6 the cascaded PI controllers achieve a success rate of $100\%$ for all types of platform movement. However, the advantage of our learning based method is that it does not require a tuning procedure as is often conducted manually for cascaded PI controllers and therefore time consuming. This is compensated by a smoother flight behaviour enabled by the use of a continuous value range for the control effort of the PI controller.

### v3

@@ 1 Introduction
Sect. 2 presents related work, followed by Sect. 3 describing the proposed approach in detail. Sect. 4 presents the implementation and experimental setup. The results are discussed in Sect. 5, with comments on future work given in Sect.  6.
@@ 3.3.1 Dynamic Decoupling
The nonlinear dynamics of a multi-rotor vehicle such as a quadcopter can be greatly simplified when the following assumptions are applied. i) Linearization around hover flight, ii) small angles iii) a rigid, symmetric vehicle body, iv) no aerodynamic effects [26]. Under these conditions, the axes of motion are decoupled. Thus, four individual controllers can be used to independently control the movement in longitudinal, lateral, vertical direction and around the vertical axis. For this purpose, low-level controllers track a setpoint $\theta_{ref}$ for the pitch angle (longitudinal motion), $\phi_{ref}$ for the roll angle (lateral motion), $T_{ref}$ for the total thrust (vertical motion) and $\psi_{rel}$ for the yaw angle (rotation around vertical axis). We leverage this fact in our controller structure as is further illustrated in Sect. 4. Furthermore, this enables us to introduce a learning task for longitudinal 1D motion only. Its purpose is to learn a policy for the pitch angle setpoint $\theta_{ref}$ that induces a 1D movement in longitudinal direction, allowing the vehicle to center over the platform moving in longitudinal direction as well. After the training, we then apply a second instance of the trained agent for controlling the lateral motion, where it produces set point values for the roll angle $\phi_{ref}$. This way, full 2D motion capability is achieved. We use the basis learning task to compose a curriculum later in Sect. 3.4. PID controllers ensure that $\phi_{rel}=0$\mathrm{r}\mathrm{a}\mathrm{d}$,\psi_{rel}=0$\mathrm{r}\mathrm{a}\mathrm{d}$$ and $v_{z_{rel}}=\text{const}<0$\mathrm{m}\mathrm{/}\mathrm{s}$$ during training.
@@ 3.3.4 Discrete State Space
where $p_{max},v_{max},a_{max}$ are the values used for the normalization of the observations. The reason for the clipping is that our state space discretization technique, a crucial part of the sequential curriculum, is derived assuming a worst case scenario for the platform movement in which the multi-rotor vehicle is hovering (see Sect. 3.4) and where scaling an observation with its maximum values, $p_{max},v_{max}$ and $a_{max}$, would constitute a normalization. However, once the multi-rotor vehicle starts moving too, the scaled observation values could exceed the value range of $[-1,1]$. Clipping allows the application of the discretization technique also for a moving multi-rotor vehicle. Next, we define a general discretization function $d(x,x_{1},x_{2})$ that can be used to map a continuous observation of the environment to a discrete state value.
@@ 3.3.5 Goal State
$s^{*}=\left\{1,1,*,*\right\}.$ (18)
This means the goal state is reached if $-p_{goal}\leq p_{x}\leq p_{goal}$ and $-v_{goal}\leq v_{x}\leq v_{goal}$ regardless of the values of $a_{d}$ and $i_{\theta}$.
@@ 3.3.6 Reward Function
The weights are negative so that decreasing the relative position, relative velocity or the pitch angle yields positive reward values. Clipping $r_{p}$ and $r_{v}$ to ±$\pm r_{p,max}$ and ±$\pm r_{v,max}$, as well as scaling $r_{\theta}$ and $r_{dur}$ with $v_{lim}$ is necessary due to their influence on the value of $r_{max}$. $r_{max}$ denotes the maximum achievable reward in a non-terminal timestep if $v_{x}\leq v_{lim}$ and $a_{x}\leq a_{lim}$ and thus complies with the limits set by the motion scenario described in Sect. 3.4 during the derivation of the sequential curriculum. To ensure the applicability of the curriculum also in situations in which that compliance is not given is why the clipping is applied in (20) and (21). $r_{max}$ plays a role in the scaling of Q-values that is performed as part of the knowledge transfer during different curriculum steps. It is defined as follows
@@ 3.4 Curriculum and Discretization
In our novel approach, we introduce a similar multiresolution technique combined with transfer learning as part of a sequential curriculum [17] to accelerate and stabilize training even further. Each new round of training in the multiresolution setting constitutes a different curriculum step. A curriculum step is an instance of the basis learning task introduced in Sect. 3.3. This means in particular that each curriculum step has its own Q-table. Since the size of the state and action space is the same throughout all steps, knowledge can be easily transferred to the subsequent curriculum step by copying the Q-table that serves then as a starting point for its training. Fig. 3 illustrates the procedure of the sequential curriculum.
@@
Furthermore, the discretization introduced in Sect. 3.3.4, which is now associated with a curriculum step, covers only a small part of the continuous observation space with few discrete states. In order to completely capture the predominant environment characteristics of the subspace in which the agent is expected to learn the task, it is necessary for it to have the knowledge of the state space topology [1]. Against the background of the multiresolution approach, this can be rephrased as the question of how to define suitable goal regions, i.e., the size of the goal states. To this end, we again leverage the insight that accurate knowledge about the relative position and velocity is only required when landing is imminent. We begin by choosing an exponential contraction of the size of the unnormalized goal state $p_{c,x,goal,i}$ associated with the $i^{th}$ curriculum step such that
@@
where $10<\sigma<$ 1 is a contraction factor, $l_{mp}\in\mathbb{R}[$\mathrm{m}$]$ is the edge length of the squared moving platform, $x_{max}$ is the boundary of the fly zone $\mathcal{F}$ (see Sect. 3.1.2) and $n_{cs}$ is the number of curriculum steps following the initial step. The contraction factor $\sigma$ plays a role in how easy it is for the agent to reach the goal state. If it is set too low, the agent will receive success rewards only rarely, thus leading to an increased training time. (34) can be solved for $n_{cs}$ because $\sigma$, $l_{mp}$ and $x_{max}$ are known parameters. To determine the relationship between the contraction of the positional goal state and the contraction of the velocity goal state, we need a kinematics model that relates these variables. For this purpose, we envision the following worst case scenario for the relative motion (1). Assume that the vehicle hovers right above the platform which is located on the ground in the center of the fly zone $\mathcal{F}$, with $v_{x}=0$\mathrm{m}\mathrm{/}\mathrm{s}$$ and $\psi_{rel}=0$. Now, the platform starts to constantly accelerate with the maximum possible acceleration $a_{mp,max}$ defined by (7) until it has reached $x_{max}$ after a time $t_{0}=\sqrt{2x_{max}/a_{mp,max}}$. This is considered to be the worst case since the platform never slows down, unlike the rectilinear periodic movement (6) used for training. The evolution of the continuous observations over time is then expressed by
@@
which constitute a simple kinematics model that relates the relative position and relative velocity via the time $t$. Next, we consider $p_{max}=x_{max}$, $v_{max}=a_{mp,max}t_{0}$ and $a_{max}=a_{mp,max}$ for the normalization (11) - (13) and discretize (35) - (37) via the time $t_{i}=\sigma^{i}t_{0}$. This allows us to determine the values of $p_{lim,i}$, $v_{lim,i}$, $a_{lim,i}$ and $p_{goal,i}$, $v_{goal,i}$, $a_{goal,i}$. They are required for the basis learning task presented in Sect. 3.3 that now constitutes the $i^{th}$ curriculum step. For this purpose, we define with $i\in\left\{0,1,\ldots,n_{cs}\right\}$
@@
During the training of the different curriculum steps the following episodic terminal criteria have been applied. On the one hand, the episode terminates with success and the success reward $r_{suc}$ is received if the goal state $s^{*}$ of the latest curriculum step is reached if the agent has been in that curriculum step’s discrete states for at least one second without interruption. This is different to all previous curriculum steps where the success reward $r_{suc}$ is received immediately after reaching the goal state of the respective curriculum step as is explained in Sect. 3.3.6. On the other hand, the episode is terminated with a failure if the multi-rotor vehicle leaves the fly-zone $\mathcal{F}$ or if the success terminal criterion has not been met after the maximum episode duration $t_{max}$. The reward received by the agent in the terminal time step for successful or failing termination of the episode is defined by (29). Once the trained agent is deployed, the agent selects the actions based on the Q-table of the latest curriculum step to which the continuous observations can be mapped (see Fig. 3).
@@ 3.5 Hyperparameter Determination
We leverage the discrete action space (8) to determine the hyperparameters agent frequency $f_{ag}$ and maximum pitch angle $\theta_{max}$ in an interpretable way. The purpose is to ensure sufficient maneuverability of the UAV to enable it to follow the platform. For sufficient maneuverability, the UAV needs to possess two core abilities. i) Produce a maximum acceleration bigger than the one the platform is capable of and ii) change direction of acceleration quicker than the moving platform. Against the background of the assumptions explained in Sect. 3.3.1 complemented with thrust compensation, we consider the first aspect by the maximum pitch angle $\theta_{max}$
@@ 4.7 Cascaded PI Controller Baseline
We apply separate instances for controlling the longitudinal, lateral and vertical movement of the UAV. The properties of the different instances are summarized in Table 4. The outer cascade tracks the setpoint $0$\mathrm{m}$$ for the relative position ${}_{s}r_{rel}$ between the multi-rotor vehicle and the moving platform. For this purpose, it generates a setpoint for the relative velocity ${}_{s}\dot{r}_{rel}$ that is tracked by the inner cascade. The inner cascade aims to reach the velocity setpoint by computing setpoints for the attitude angle $\phi_{ref},\theta_{ref}$ or the thrust $T_{ref}$, respectively. The attitude angle or thrust is then tracked by the multi-rotor vehicle’s low level controllers. In order to achieve comparability between our RL approach and the cascaded PI controller, we limit the velocity setpoints generated by the outer cascade to the value of $v_{max}=a_{mp,max}t_{0}$ (see Sect. 3.4) and the attitude angles generated by the inner cascade to the same angles determined using (44) for the training scenario associated with the agent chosen for comparison.
@@ 3 $9,7,8,8$ $8,6,7,7$ $7,6,8,8$ $6,7,6,8$
They were determined over a larger set of RPMs than in the baseline method and indicate a good performance of the approach. The success rates become higher when the platform velocity is lower during evaluation compared to the one applied during training. This is to be expected since the equations presented in Sec. 3.5 for hyperparameter determination ensure a sufficient maneuverability up to the velocity of the rectilinear periodic movement used in training. For the training case RPM 0.8, the fourth agent has a comparably poor performance. For the static platform the agent occasionally suffers from the problem of motion initiation as described in Sect. 4.6, despite setting $\psi_{rel}=\pi/4$\mathrm{r}\mathrm{a}\mathrm{d}$$. For the evaluation case where the platform moves with RPM 0.4, the reason is an oscillating movement, which causes the agent to occasionally overshoot the platform. A similar problem arises for agent four of training case hardware where the agent occasionally expects a maneuver of the platform during the landing procedure. As a consequence, this agent achieves a success rate of only $82\%$ for a static platform. However, for higher platform velocities, the success rate improves significantly.
@@ 5.2 Selection of Agents for further Evaluation
We compare its results with the RL baseline method and the cascaded PI controller in Table 8. The comparison shows that our approach is able to outperform the RL baseline method. For the RPM $0.4$ evaluation scenario with noiseless observations we achieve a success rate of $99\%$, which is $+8\%$ better than the baseline. For the RPM $1.2$ evaluation scenario, our method is successful in $99\%$ of the landing trials, increasing the RL baseline’s success rate by $+26\%$. Note that the maximum radius of the rectilinear periodic movement is $r_{mp}=\sim 2.5$\mathrm{m}$$ in the RL baseline and $r_{mp}=2$\mathrm{m}$$ in our approach. However, the value used for our approach poses a more difficult challenge since the acceleration acting on the platform is higher, due to the same maximum platform velocity, see Table 2. Furthermore, our method requires $\sim 80\%$ less time to train and $53\%$ less episodes. Comparison with the cascaded PI controller shows that our method is capable of similar performance than the cascaded PI controller in terms of success rates achieved in the two comparison scenarios. In order to achieve comparability in the first place, we limited the control efforts of the longitudinal and lateral controller’s outer and inner cascade (see Table 4) to values also derived for the training case RPM 1.6. We obtain $u_{o}=v_{max}=a_{mp,max}t_{0}=3.39$\mathrm{m}\mathrm{/}\mathrm{s}$$ as described in Sect. 3.4 and $u_{i}=\theta_{ref}=\phi_{ref}=21.38^{\circ}$ by means of (44), respectively. As is further shown by Table 6 the cascaded PI controllers achieve a success rate of $100\%$ for all types of platform movement. However, the advantage of our learning based method is that it does not require a tuning procedure as is often conducted manually for cascaded PI controllers and therefore time consuming. This is compensated by a smoother flight behaviour enabled by the use of a continuous value range for the control effort of the PI controller.